Stubbs Company uses the perpetual inventory method. On January 1, 2016, Stubbs purchased 400 units of inventory that cost $8.00 each. On January 10, 2016, the company purchased an additional 600 units of inventory that cost $9.00 each. If Stubbs uses a weighted average cost flow method and sells 700 units of inventory for $16.00 each, the amount of gross profit reported on the income statement will be:

A. $5,180. B. $5,250. C. $5,000. D. $6,020.
Answer #1

Total cost = (400*8)+(600*9) = 8600

Total units = 400 + 600 = 1000

Average cost per unit = 8600/1000 = 8.60

Gross profit = Sales revenue - Cost of Goods Sold

= (700*16) - (700*8.60)

= 11,200 - 6020

= 5180